OpenStudy (anonymous):

the recipe for a loaf cake calls for 4/5 pound of jelly beans in assorted colors. If 25 cakes are needed, how many pounds of jelly beans will be used? ( simplify your answer. Type an integer or a fraction)

OpenStudy (anonymous):

1 cake = 4/5 lbs of jelly beans 25 cakes = ??? lbs of jelly beans I would say : how do you go from 1 cake to 25 cakes: Multiply by 25 so I know that both sides must do the same thing so I say, 4/5 * 25 = 20 lbs of jelly beans needed

OpenStudy (anonymous):

so would i type 20 lbs as the answer

OpenStudy (anonymous):

Yes, but do you understand how I did that?

OpenStudy (anonymous):

yes you explained it well

OpenStudy (anonymous):

Great!

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

uw